OpenStudy (anonymous):

Suppose that y varies directly as x, and y = 21 when x = 9. What is the value of y when x = –6?

OpenStudy (lgbasallote):

direct variation so y = kx when y = 21 x = 9 21 = 9k k = 21/9 what is y when x = -6 y = kx y = (21/9)(-6) can you do this now?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

you just have to use a calculator o.O

OpenStudy (lgbasallote):

you know how to use a calculator right?

OpenStudy (anonymous):

just tell me if y = -14

OpenStudy (lgbasallote):

it is